Generate a general formula, in simplest form, that will give the maximum volume of a rectangular based prism made from a rectangular piece of paper. The dimensions of the rectangular page are w ×2w with a square of dimensions x×x cut from each corner.
1
Expert's answer
2019-07-31T11:08:34-0400

As per the question,

the length will be "2w-2x,"

width will be "w-2x,"

height will be "x"

of the rectangular prism.

The required volume will be "x(2w-2x)(w-2x)"

So,

"V=x(2w-2x)(w-2x)"

"V=2w^2x-6wx^2+4x^3"

For V to be maximum,"\frac{dV}{dx}=0"


"\frac{dV}{dx}=2w^2-12wx+12x^2=0"


Solving this equation,

we get,

"6x^2-6wx+w^2=0"

By using quadratic formula


"x=\frac{-b+\sqrt{b^2-4ac}}{2a}"

and


"x=\frac{-b-\sqrt{b^2-4ac}}{2a}"


Discriminant(D)="\sqrt{b^2-4ac}"

So

"x=\frac{6w-\sqrt{36w^2-24w^2}}{12}"

(Neglected positive discriminant because "w-2x" would be negative for that case, which is not possible)


"x=\frac{(3-\sqrt{3})w}{6}"

which can be further written as

"x=\frac{w}{3+\sqrt{3}}"

So by using value of "\sqrt{3}=1.732"

we can solve and get,

"x=0.211w"